40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Wed Jul 15 22:23:17 2015, Koji, Electronics, AM Stabilized EOM Driver, E1400445 first look IMG_20150714_195534852.jpgIMG_20150714_195227746_HDR.jpg
    Reply  Sat Jul 18 11:37:21 2015, Koji, Electronics, AM Stabilized EOM Driver, D0900848 power board ~ oscillation issue solved 6x
       Reply  Wed Jul 22 09:43:01 2015, Koji, Electronics, AM Stabilized EOM Driver, Power supply test of the EOM/AOM Driver 
          Reply  Wed Jul 22 10:15:14 2015, Koji, Electronics, AM Stabilized EOM Driver, RF test of the EOM/AOM Driver S1500117 
             Reply  Sat Jul 25 17:24:11 2015, Koji, Electronics, AM Stabilized EOM Driver, RF test of the EOM/AOM Driver S1500117 7x
                Reply  Tue Jul 28 18:36:50 2015, Koji, Electronics, AM Stabilized EOM Driver, RF test of the EOM/AOM Driver S1500117 IMG_20150727_214536773_HDR.jpgEOM_Driver_DAQ_TF_test.pdfEOM_Driver_Mon_PSD.pdf
Message ID: 227     Entry time: Wed Jul 22 09:43:01 2015     In reply to: 225     Reply to this: 228
Author: Koji 
Type: Electronics 
Category: AM Stabilized EOM Driver 
Subject: Power supply test of the EOM/AOM Driver 

Serial Number of the unit: S1500117
Tester: Koji Arai
Test Date: Jul 22, 2015

1) Verify the proper current draw with the output switch off:

+24 Volt current: 0.08 A Nom.
-24 Volt current: 0.07 A Nom.
+18 Volt current: 0.29 A Nom.
-18 Volt current: 0.24 A Nom.

2) Verify the proper current draw with the output switch on:
+24 Volt current: 0.53 A Nom.
-24 Volt current: 0.07 A Nom.
+18 Volt current: 0.21 A Nom.
-18  Volt current: 0.26 A Nom.

3) Verify the internal supply voltages:

All look good.

TP13 -5.001V
TP12 -15.00
TP11 -21.05
TP10 -10.00
TP5  -18.19
TP6  -24.22
TP2  +24.15
TP3  +18.22
TP9  + 9.99
TP17 +24.15
TP14 +21.04
TP15 +15.00
TP16 + 4.998

4) Verify supply OK logic:
All look good. This required re-disassembling of the PCBs...

Check then pin 5 on U1 (connected to R11) and U4 (connected to R23):

U1 3.68V (=Logic high)
U4 3.68V (=Logic high)

5) Verify the relays for the power supply sequencing: OK

Turn off +/-24 V. Confirm Pin 5 of K1 and K2 are not energized to +/-18V. => OK
Turn on +/-24 V again. Confirm Pin 5 of K1 and K2 are now energized to +/-18V. => OK

6) Verify noise levels of the internal power supply voltages:

TP13 (- 5V) 13nV/rtHz@140Hz
TP12 (-15V) 22
nV/rtHz@140Hz
TP11 (-21V) 32
nV/rtHz@140Hz
TP10 (-10V) 16nV/rtHz@140Hz
TP9  (+10V)  9
nV/rtHz@140Hz
TP14 (+21V) 21nV/rtHz@140Hz
TP15 (+15V) 13nV/rtHz@140Hz
TP16 (+ 5V) 11nV/rtHz@140Hz

 

Note that the input noise of SR785 is 9~10nV/rtHz@140Hz with -50dBbpk input (AC)

 

ELOG V3.1.3-